OpenStudy (anonymous):

Evaluate the following expression for r = -5 and t = 8 5r + 9t – 3rt + 4 HELP PLEAS

OpenStudy (jack1):

5r + 9t – 3rt + 4 = 5x-5 + 9x8 - 3x-5x8 + 4

OpenStudy (jack1):

= -25 + 72 + 4 + 120

OpenStudy (jack1):

following ok?

OpenStudy (anonymous):

5r + 9t – 3rt + 4 5x(-5)+9(8)-3(-5x8)+4 -25+72+40+4 91

OpenStudy (anonymous):

No you forgot 3rt :/

OpenStudy (jack1):

how is 3 x 5 x 8 = 40? @xoya

OpenStudy (anonymous):

sorry right :(

OpenStudy (anonymous):

91 is not one of my answer choices @xoya but thanks for the help

OpenStudy (anonymous):

5r + 9t – 3rt + 4 5x(-5)+9(8)-3(-5x8)+4 -25+72+120+4 171

OpenStudy (jack1):

@ur-only_angel add these: = -25 + 72 + 4 + 120

OpenStudy (anonymous):

I did it :) now its correct :)

OpenStudy (anonymous):

thank you both @xoya and @jake1(:

OpenStudy (jack1):

yw

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

@ur-only_angel 171 you have in choices? :)